Circular No. 6701

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1996 H1 (SOHO) C. St. Cyr, Naval Research Laboratory, reports, for the SOHO-LASCO

Consortium (see IAUC 6685), the discovery by B. McCarty and D. Biesecker of another Kreutz sungrazer in SOHO-LASCO C3 coronagraph archival data. Measurements by Biesecker have been reduced by G. V. Williams and are given on MPEC 1997-O05 in detail, the approximate discovery position being:

1996 UT  α_{2000}  δ_{2000}
Apr. 29.155  2^{h}43.^{m}6  +11^{o}52′

The comet was estimated to vary between mag 5 and 9, and there may have been a tail in the single C2 image (on Apr. 30.202 UT). The orbit computed by the undersigned, also given in detail on MPEC 1997-O05, indicates perhelion passage on Apr. 30.50.

η CARINAE

M. F. Corcoran, Goddard Space Flight Center (GSFC) and Universities Space Research Association; J. H. Swank and R. Petre, GSFC; and K. Ishibashi and K. Davidson, University of Minnesota, report: “Target of Opportunity observations with the Rossi X-ray Timing Explorer (RXTE) of the peculiar luminous star η Car starting June 20 show significant x-ray variability on timescales of days and confirm the presence of an x-ray peak on July 4.96 UT. This is close to the time of the predicted peak based on the 85.1-day period previously reported (IAUC 6668), and it strongly suggests that at least some of the 2–10-keV flux variability is periodic. Using a conversion factor of 3.22×10^{−12} erg s^{−1} cm^{−2} per PCA count, the backgroundcorrected flux in the 2–10-keV band was 1.29 × 10^{−10} erg s^{−1} cm^{−2} on July 4.96. This was 7.5 percent larger than the previous peak on Apr. 10. The 2–10-keV flux declined to 1.05×10^{−10} erg s^{−1} cm^{−2} on July 11. Continued monitoring at x-ray and other energies is needed to identify the cause of the periodic emission. Dates of upcoming and previous RXTE observations of η Car, as well as recent RXTE lightcurves, are available from http://

lheawww.gsfc.nasa.gov/users/corcoran/eta car/eta car xte.html.”

SUPERNOVA 1997cx IN NGC 3057 S. Jha, Harvard-Smithsonian Center for Astrophysics, notes that the

Whipple Observatory images of SN 1997cx (cf. IAUC 6700) show it to be 5′′ west and 15′′ north (not south) of the center of NGC 3057.

1997 July 17  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Brian G. Marsden